OpenStudy (anonymous):

Solve the equation and check for extraneous equations. 9|9-8x|=2x+3

OpenStudy (anonymous):

\[ |9-8x| = \frac{2x+3}{8} \]\[ \begin{array}{} -(9-8x) = \frac{2x+3}{8} &\text{and} & +(9-8x) = \frac{2x+3}{8} \end{array} \]Solve them both individually.

OpenStudy (anonymous):

Is the first one x=12.5 and the other is x=6.9? im still confused on whats the answer..